# German Research Center for Artificial Intelligence (DFKI), Saarbrücken, Germany

### v1

@@ 2 Related Work
In the domain of multilingual NLP, models like mT5[11] have demonstrated strong performance across multiple languages by pre-training on a massive multilingual corpus. Similarly, several works have focused on multilingual IR. ColBERT-X[8] extends the ColBERT[6] architecture to multilingual settings, showing effectiveness in languages included in its training data. Benchmarks like MIRACL[12] and initiatives like CIRAL[1] aim to standardize evaluation and improve cross-lingual retrieval. However, these works often do not encompass low-resource languages like Urdu.
@@ 3.1 Ms-Marco dataset
The MS MARCO Passage ranking dataset [3]  [^2] https://microsoft.github.io/msmarco/Datasets.html is a cornerstone in IR research, comprising over 8.8 million passages and more than 500 thousand unique queries each paired with at least one relevant passage. The commonly used development set includes 6,980 queries, while the training set is organized into 39 million triples, each consisting of a query, a relevant passage, and a non-relevant passage. An example of it can be seen in Table 1. This dataset’s scale and richness have significantly advanced English language IR models. However, replicating such progress for Urdu is challenging due to the absence of equivalent large-scale datasets or the prohibitive costs associated with translating such extensive corpora333estimated around \$30,000 for MS Marco, using services like Azure Translator or Google Translation API with their standard rates of 10-20\$ per million characters.
@@ 3.2 Translation Process
To translate the dataset, we used the open-source IndicTrans2[5] model. We utilized the distilled 200-million parameter version of IndicTrans2 to translate the entire MS MARCO dataset into Urdu. The process involved tokenizing the text with the IndicTransTokenizer, translating in batches to optimize computational efficiency and leveraging GPU acceleration. Key parameters included a batch size of 32 and a maximum sequence length of 512 tokens to handle lengthy passages. We also performed preprocessing steps, such as normalization and error correction to enhance the quality and usability of the dataset and the model. Overall the translation took around 120 hours to translate the dataset on a single V100 GPU with 32GB Vram.
Table 1: Example of a query and its relevant and non-relevant passage in both English and machine-translated Urdu
@@ 4.3 Evaluation Metrics
To assess the effectiveness of our models, we adopted the same evaluation metrics as the original mMarco benchmark, ensuring comparability. The Mean Reciprocal Rank at 10 (MRR@10) was calculated to evaluate the average reciprocal rank of the first relevant passage retrieved within the top 10 results for each query. This metric provides insight into the ranking quality of the retrieval system, emphasizing the position of the first relevant result. Additionally, we computed Recall at 10 (Recall@10) to measure the proportion of relevant passages retrieved in the top 10 results out of all relevant passages available for a query. This metric reflects the system’s ability to retrieve as many relevant documents as possible within a limited set of results. Together, these metrics offer a balanced evaluation of both the precision and recall aspects of the retrieval performance.
@@
Model Dataset Language Recall@10 MRR@10
BM25 (k=1000) English 0.391 0.187
mMARCO Reranker English - 0.370
BM25 (k=1000) Urdu 0.268 0.129
Zero-Shot mMARCO Reranker Urdu 0.408 0.204
Urdu mT5-mMARCO ( This Work ) Urdu 0.438 0.247
@@ 5.1 Performance Comparison
We evaluated the retrieval performance on the Urdu dataset using BM25 for initial retrieval, followed by different re-ranking strategies (see Table 2). The baseline BM25 model is used to retrieve the top 1000 passages per query and it achieves an MRR@10 of 0.129 on the Urdu dataset. This performance is significantly lower than BM25 results on the English dataset and is among the lowest scores reported for other languages in prior studies [4]. This highlights the challenges of applying traditional retrieval methods to low-resource languages like Urdu, likely due to morphological complexity and limited linguistic resources.
@@
Our fine-tuned mMARCO model, trained on the Urdu-translated dataset, provided further improvements when used as a re-ranker over BM25 outputs. It achieved the highest performance with an MRR@10 of 0.247 and a Recall@10 of 0.438. Fine-tuning allowed the model to adapt to the linguistic characteristics of Urdu, enhancing its ability to distinguish relevant passages more effectively than the zero-shot model. These gains demonstrate that language-specific fine-tuning, even on machine-translated data, can significantly boost retrieval effectiveness in low-resource languages.
@@ 6 Future Work and Conclusion
This work contributes to more inclusive IR by making retrieval systems accessible to Urdu speakers, with implications for other South Asian languages. Future work could involve incorporating manual verification to refine translation quality and exploring hybrid methods that combine machine and human translations. Expanding this approach to other low-resource languages would further support the development of fair, balanced and multilingual IR technologies.

### v3

# German Research Center for Artificial Intelligence (DFKI), Germany

@@ 1 Introduction
  • Publicly Available Resources: To facilitate future research in Urdu IR, we make our translated dataset, fine-tuned mT5 model, and code publicly available on Hugging Face  [^1] https://huggingface.co/datasets/Mavkif/urdu-msmarco-dataset  [^2] https://huggingface.co/Mavkif/urdu-mt5-mmarco and GitHub  [^3] https://github.com/UmerTariq1/Urdu_MsMarco_Translation_Retrieval.
@@ 2 Related Work
In the domain of multilingual NLP, models like mT5[11] have demonstrated strong performance across multiple languages by pre-training on a massive multilingual corpus. Similarly, several works have focused on multilingual IR. ColBERT-X[8] extends the ColBERT[6] architecture to multilingual settings, showing effectiveness in languages included in its training data. Benchmarks like MIRACL[12] and initiatives like CIRAL[1] aim to standardize evaluation and improve cross-lingual retrieval, particularly for low-resource scenarios. MIRACL addresses this by providing a diverse set of 18 languages, including both high and low-resource languages from 10 different language families, with a focus on monolingual retrieval. It features a variety of tasks, including question answering and document retrieval, and incorporates challenges like code-switching and dialectal variations. While MIRACL includes a diverse set of languages, it currently lacks a dedicated Urdu corpus.
@@ 3.1 Ms-Marco dataset
The MS MARCO Passage ranking dataset [3]  [^5] https://microsoft.github.io/msmarco/Datasets.html is a cornerstone in IR research, comprising over 8.8 million passages and more than 500 thousand unique queries each paired with at least one relevant passage. The commonly used development set includes 6,980 queries, while the training set is organized into 39 million triples, each consisting of a query, a relevant passage, and a non-relevant passage. An example of it can be seen in Table 1. This dataset’s scale and richness have significantly advanced English language IR models. However, replicating such progress for Urdu is challenging due to the absence of equivalent large-scale datasets or the prohibitive costs associated with translating such extensive corpora666estimated around \$30,000 for MS Marco, using services like Azure Translator or Google Translation API with their standard rates of 10-20\$ per million characters.
@@ 3.2 Translation Process
To create the Urdu MS MARCO dataset, we translated the entire English MS MARCO dataset, including both passages and queries from the training and development sets, using the open-source IndicTrans2[5] model. We utilized the distilled 200-million parameter version of IndicTrans2 to translate the entire MS MARCO dataset into Urdu. The process involved tokenizing the text with the IndicTransTokenizer, translating in batches to optimize computational efficiency and leveraging GPU acceleration. Key parameters included a batch size of 32 and a maximum sequence length of 512 tokens to handle lengthy passages. We also performed preprocessing steps, such as normalization and error correction to enhance the quality and usability of the dataset and the model. Overall the translation took around 120 hours to translate the dataset on a single V100 GPU with 32GB Vram.
Table 1: Example of a query and its relevant and non-relevant passage in both English and machine-translated Urdu. More examples can be seen on huggingface dataset webpage
@@ 4.3 Evaluation Metrics
To assess the effectiveness of our models, we used common evaluation metrics in information retrieval research to ensure comparability with other studies. These metrics include Mean Reciprocal Rank at 10 (MRR@10), which evaluates the ranking quality by considering the position of the first relevant result within the top 10 retrieved passages. We also used Recall at 10 (Recall@10) to measure the proportion of relevant passages retrieved in the top 10 results, reflecting the system’s ability to find as many relevant passages as possible. These two metrics were also used in the original mMarco benchmark. Additionally, we included Mean Average Precision at 10 (MAP@10) to assess the overall ranking quality across all relevant passages for each query. Finally, we used Normalized Discounted Cumulative Gain at 10 (NDCG@10), a more sophisticated metric that takes into account the position of all relevant passages in the ranking and gives higher weight to those appearing earlier in the list. By employing these metrics, we aimed to provide a comprehensive evaluation of retrieval performance, encompassing various aspects of ranking quality and relevance.
@@
Model Dataset Language Recall@10 MRR@10 MAP@10 NDCG@10 BM25 (k=1000) English 0.391 0.187 0.104 0.275 mMARCO Reranker English 0.639 0.370 0.245 0.498 BM25 (k=1000) Urdu 0.247 0.121 0.070 0.177 Zero-Shot mMARCO Reranker Urdu 0.408 0.204 - - Urdu mT5-mMARCO ( This Work ) Urdu 0.438 0.248 0.159 0.340
@@ 5.1 Performance Comparison
We evaluated the retrieval performance on the Urdu dataset using BM25 for initial retrieval, followed by different re-ranking strategies (see Table 2). The baseline BM25 model is used to retrieve the top 1000 passages per query and it achieves an MRR@10 of 0.121 on the Urdu dataset. The baseline BM25 model exhibited significantly lower performance on the Urdu dataset compared to the English dataset. This disparity underscores the challenges of applying traditional retrieval techniques to morphologically rich languages like Urdu, where effective tokenization requires more nuanced approaches than those typically used for English. The limitations of standard tokenizers for handling Urdu morphology likely contribute to the lower performance observed with BM25.
@@
Our fine-tuned model, Urdu mT5-mMARCO, yielded the most substantial gains. Fine-tuning on the Urdu dataset allowed the model to adapt to the nuances of the language, leading to improvements across all reported metrics. Notably, Urdu mT5-mMARCO achieved a Recall@10 of 0.438, surpassing both the BM25 baseline (0.247) and the zero-shot mMARCO (0.408). Furthermore, the fine-tuned model exhibited improvements in MRR@10 (0.248), exceeding the zero-shot model (0.204) and significantly outperforming the BM25 baseline (0.121). The gains observed in MAP@10 (0.159) and NDCG@10 (0.340) further demonstrate the positive impact of fine-tuning. This highlights the effectiveness of language-specific fine-tuning, even with machine-translated training data, in enhancing retrieval performance for low-resource languages. The results demonstrate the potential of this approach to bridge the gap in information access for Urdu speakers.
@@ 6 Future Work and Conclusion
This work contributes to more inclusive IR by making retrieval systems accessible to Urdu speakers, with implications for other South Asian languages. Future work could involve incorporating manual verification to refine translation quality and exploring hybrid methods that combine machine and human translations. To better understand the impact of these methods, a comprehensive analysis of translation quality and its quantitative effect on retrieval performance will be conducted. Expanding this approach to other low-resource languages would further support the development of fair, balanced and multilingual IR technologies.
7 Acknowledgement
This research was conducted at the German Research Center for Artificial Intelligence (DFKI) and the University of Saarland. The authors gratefully acknowledge DFKI for providing the necessary hardware resources and the supportive research environment. This work was also supported by the German Federal Ministry of Education and Research (BMBF) as part of the TRAILS project (01IW24005). The authors thank the BMBF for their commitment to advancing research in Artificial Intelligence.